(2) According to the general definition of aromaticity and the Huckels 4n T rule, predict the aromaticity of the following compounds. 2. Write the numbers of aromatic compounds: Aromatic compounds are:
Answer #1

There are four aromatic compounds present.

They are options 1,2,4, and 5.

option 1 has 5 double bonds = 10 pi electrons ( 4*2 + 2 = 10 so it obeys 4n+2 rule)

option 2 has 2 double bonds and in O there are two lone pair of electrons among two lone one pair involves delocalization pi electrons hence total 6 electrons in cyclic delocalization ( 4*1 +2 = 6)

option 4 has 2 pi bond ( 4 pi electron ) and one lone pair electron on carbon with negative charge so total 6 electron

option 5 has 7 double bond = 14 electron ( 4*3 +2 = 14)